OpenStudy (anonymous):

]Henry teaches gymnastics. He charges a one−time processing fee of $15.00 and the cost of each class is shown below. Let x represent the number of classes and y represent the cost of classes. Based on this information, what will it cost, in dollars, to take 5 classes?

OpenStudy (campbell_st):

Total cost = 15 +5y general solution is TC = 15 +xy

OpenStudy (anonymous):

SO ITS 15

OpenStudy (anonymous):